# Topological and nontopological degeneracies in generalized string-net models

### v3

@@ C.2 Stability inequality
is described by a rectangular matrix $n_{A,s}$, called a restriction or lifting matrix, with $A\in\mathcal{Z}(\mathcal{C})$ and $s\in\mathcal{U}$ (in our case, $s=1$ only). Here, this matrix is the vector $\mathbf{n}_{1}$ and the condensing bosons are the fluxons. The condensation equations (see, e.g., Eqs. (20a) and (20b) in Ref. [42]) relate the $S$- and $T$- matrices of the two UMTCs $\mathcal{A}$ and $\mathcal{U}$ as
@@
As $T_{\mathcal{U}}=1$ and $S_{\mathcal{U}}=1$ are trivial $1\times 1$ matrices, and $T_{\mathcal{A}}=T$ and $S_{\mathcal{A}}=S$ are the $T$- and $S$- matrices of $\mathcal{Z}(\mathcal{C})$, we find Eq. (11-12). Anyon condensation should also hold in the case in which the UFC $\mathcal{T}$ has noncommutative fusion rules, as briefly discussed by Bais and Slingerland [56].
@@
An interesting question to ask is, given an achiral topological order characterized by a Drinfeld center $\mathcal{Z}(\mathcal{C})$ with modular matrices $S$ and $T$, what are the possible gapped boundaries or condensations to the vacuum [42, 44, 57] ?
A partial answer to that question is known (see, e.g., Refs [44, 58, 57]): a necessary (but not sufficient) condition for a gapped boundary is to have a vector $\mathbf{n}_{1}$ that satisfies the fluxon identities Eqs. (11)-(12) and the stability condition (98), i.e., to have a stable fluxon set. This is equivalent to the notion of Lagrangian algebras in $\mathcal{Z}(\mathcal{C})$ [59]. For each such stable $\mathbf{n}_{1}$, there is a corresponding boundary theory described by a UFC $\mathcal{C}_{b}$. Obviously, $\mathcal{C}$ is one of the possible boundary theories $\mathcal{C}_{b}$. Also, all the boundary theories are Morita-equivalent, i.e. $\mathcal{Z}(\mathcal{C}_{b})\simeq\mathcal{Z}(\mathcal{C})$. This is the bulk-boundary correspondence for achiral topologically ordered phases.
For example, given $\mathcal{Z}(S_{3})$ (see Sec. VI.1 for the notations), one finds five possible $\mathbf{n}_{1}$’s satisfying (11) and (12), one of which [$\mathbf{n}_{1}=\{1,1,1,0,0,1,0,0\}^{T}$ which for short we note $\mathbf{n}_{1}=(A,B,C,F)$ according to its non-zero elements] does not satisfy the stability condition (98[57]. Among the four stable solutions, two are related by symmetry and correspond to Vec($S_{3}$) [i.e. $\mathbf{n}_{1}=(A,B,2\times C)$ and $\mathbf{n}_{1}=(A,B,2\times F)$], and two are related by symmetry and correspond to Rep($S_{3}$) [i.e. $\mathbf{n}_{1}=(A,D,F)$ and $\mathbf{n}_{1}=(A,C,D)$], accounting for the symmetry between the $C$ and $F$ quasiparticles [49].
Another example is that of $\mathcal{Z}(\mathcal{H}_{3})$ (see Sec. VI.3 for the notations). In this case, we find three stable solutions. Two solutions related by symmetry [i.e. $\mathbf{n}_{1}=\{1,0,0,0,0,0,0,1,2,0,0,0\}^{T}=(\mathbf{1},\pi_{1},2\times\pi_{2})$ and $\mathbf{n}_{1}=(\mathbf{1},\pi_{1},2\times\sigma_{1})$], due to the symmetry between the $\pi_{2}$ and $\sigma_{1}$ quasiparticles, point to the fusion ring $H_{6}$, and one [i.e. $\mathbf{n}_{1}=(\mathbf{1},\pi_{1},\pi_{2},\sigma_{1}$)] points to the fusion ring $H_{4}$ [60]. The fusion ring $H_{6}$ corresponds to the UFCs $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$, whereas $H_{4}$ refers to the UFC $\mathcal{H}_{1}$. The two categories $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$ have the same fusion rules but different $F$ symbols.
@@ Appendix D Generalized Hamiltonian
The Hamiltonian that we consider in the main text, Eq. (10), assigns the same energy penalty +1 to each fluxon that is not the vacuum. However, we can assign different energy penalties to each fluxon type. Further, we can also assign a plaquette-dependent energy penalties. More generally, let us define a projector
@@
where $E^{A}_{p}$ is the energy cost of having fluxon $A$ on plaquette $p$, and $\cal F$ is the set of fluxons. As long as $E^{\bf 1}_{p}<E^{A}_{p}$ for all $A\neq{\bf 1}$ and all $p$, then the ground state is still the state where all plaquettes are in the vacuum state. The Hamiltonian we consider in the main text corresponds to $E_{p}^{A}=0$ for all $A\neq{\bf 1}$ and $E_{p}^{\bf 1}=-1$, for all plaquettes $p$.
It is not hard to construct the projectors $B_{p}^{A}$ from the structure of the tube algebra by inserting a projector $P_{A}^{11}$ inside the plaquette $p$ and then fusing it into the edges of the plaquette. For the case in which $n_{A,1}>1$, in fact there is freedom to generalize Hamiltonian Eq. (99) further. Since the space of states where a plaquette $p$ has fluxon $A$ is $n_{A,1}$-dimensional, one can add terms to split this added degeneracy. Thus we can even more generally write
@@
Let us take an example of the category Vec($S_{3}$) as discussed in Sec. VI.1. There are three fluxon types labeled $A,B,C$. Here $A$ is the vacuum, which we give energy $E^{A}=0$. Then for $B$ we give energy $E^{B}=x$. Finally $C$ is two-dimensional, i.e. $n_{C,1}=2$, and we call $E^{C,1}=y$ and $E^{C,2}=z$ the two values of $E^{C,a}$ with $a=1,2$. For simplicity here, we assume that all plaquettes are identical to each other, although this is not necessary.
@@ Appendix E Hilbert-space dimension
Both calculations are actually sums over fusion networks. In Sec. E.1, we are summing over fusion trees of $\cal C$ whereas, in Sec. E.2, we are summing over fusion trees of $\cal Z(\cal C)$ (see the discussion on the derivation of the Moore-Seiberg-Banks formula in Sec. IV.1). In both cases, the fusion trees are sums over fusion multiplicity matrices. In Sec. E.1 we are summing fusion matrices of $\cal C$ whereas in Sec. E.2 we are summing fusion matrices of ${\cal Z}({\cal C})$. These two different fusion multiplicity matrices are related to each other via the commutativity of fusion and restriction [56, 42]. In particular, we have Eq. (97), where $r,s,t\in\cal C$ and $A,B,C\in\cal Z(C)$ and $n_{A,r}$ are lifting coefficients. To clearly distinguish the two sets of fusion matrices we use $\tilde{N}$ for the input category $\mathcal{C}$ and $N$ for the output category $\mathcal{Z}(\mathcal{C})$.
@@ E.1 Counting edge labelings
For more complex graphs, one may use restructuring moves (or “$F$-moves,” which here are just the associativity of fusion) to change the connectivity of the graph without altering the Hilbert-space dimension, as shown in Fig. 14.
@@
However, in the case in which fusion rules are noncommutative, the situation is different, and on a surface of nonzero genus, one cannot reduce the fusion diagram to a chain of bubbles of the form $L$. To understand this, let us first consider the case $g=1$ (torus). We draw the torus as a rectangle with opposite edges identified. Using the above restructuring moves (Fig. 14), we can reduce any trivalent graph to the following (where the graph edges are black, and we do not draw arrows or labels on these edges for simplicity).
@@
If there are multiple lines running around the handles, Eq. (106) can be reduced to single lines running around the handle by using restructuring moves (Fig. 14). However if there is a single line going around a handle (as shown in the above diagram), it cannot be removed by restructuring moves.
@@
where the over-crossing could be drawn equally well as an undercrossing, since it is only the connectivity that matters. But note that due to noncommutativity of the fusion rules, it is not possible to remove the under-crossing and flatten the diagram.
@@
since two vertices occur within the twisted bubble, and the remaining vertices are within the planar graph, which can be reduced to a bubble chain of $L$’s. Note that for the case of commutative fusion, we have $\tilde{L}=L$ and we recover the previous result (105).
To extend this to a genus-$g$ surface, we start with the planar representation of this surface as a $4g$-sided polygon with certain edges identified as in the following picture:
@@
Unfortunately, this structure is still somewhat inconvenient, so we use restructuring moves to connect the legs of each twisted bubble, getting
@@
The total Hilbert-space dimension is then

### v4

@@ C.2 Stability inequality
is described by a rectangular matrix $n_{A,s}$, called restriction or lifting matrix, with $A\in\mathcal{Z}(\mathcal{C})$ and $s\in\mathcal{U}$ (in our case, $s=1$ only). Here, this matrix is the vector $\mathbf{n}_{1}$ and the condensing bosons are the fluxons. The condensation equations (see, e.g., Eqs. (20a) and (20b) in Ref. [42]) relate the $S$- and $T$- matrices of the two UMTCs $\mathcal{A}$ and $\mathcal{U}$ as
@@
As $T_{\mathcal{U}}=1$ and $S_{\mathcal{U}}=1$ are trivial $1\times 1$ matrices, and $T_{\mathcal{A}}=T$ and $S_{\mathcal{A}}=S$ are the $T$- and $S$- matrices of $\mathcal{Z}(\mathcal{C})$, we find Eq. (11-12). Anyon condensation should also hold in the case where the UFC $\mathcal{T}$ has noncommutative fusion rules, as briefly discussed by Bais and Slingerland [56].
@@
An interesting question to ask is: given an achiral topological order characterized by a Drinfeld center $\mathcal{Z}(\mathcal{C})$ with modular matrices $S$ and $T$, what are the possible gapped boundaries or condensations to the vacuum [42, 44, 57] ?
A partial answer to that question is known (see, e.g., Refs [44, 58, 57]): a necessary (but not sufficient) condition for a gapped boundary is to have a vector $\mathbf{n}_{1}$ that satisfies the fluxon identities Eqs. (11)-(12) and the stability condition (98), i.e., to have a stable fluxon set. This is equivalent to the notion of Lagrangian algebras in $\mathcal{Z}(\mathcal{C})$ [59]. For each such stable $\mathbf{n}_{1}$, there is a corresponding boundary theory described by a UFC $\mathcal{C}_{b}$. Obviously, $\mathcal{C}$ is one of the possible boundary theories $\mathcal{C}_{b}$. Also, all the boundary theories are Morita-equivalent, i.e. $\mathcal{Z}(\mathcal{C}_{b})\simeq\mathcal{Z}(\mathcal{C})$. This is the bulk-boundary correspondence for achiral topologically-ordered phases.
For example, given $\mathcal{Z}(S_{3})$ (see Sec. VI.1 for the notations), one finds five possible $\mathbf{n}_{1}$’s satisfying (11) and (12), one of which ($\mathbf{n}_{1}=\{1,1,1,0,0,1,0,0\}^{T}$ which for short we note $\mathbf{n}_{1}=(A,B,C,F)$ according to its non-zero elements) does not satisfy the stability condition (98[57]. Among the four stable solutions, two are related by symmetry and correspond to Vec($S_{3}$) [i.e. $\mathbf{n}_{1}=(A,B,2\times C)$ and $\mathbf{n}_{1}=(A,B,2\times F)$], and two are related by symmetry and correspond to Rep($S_{3}$) [i.e. $\mathbf{n}_{1}=(A,D,F)$ and $\mathbf{n}_{1}=(A,C,D)$], accounting for the symmetry between the $C$ and $F$ quasiparticles [49].
Another example is that of $\mathcal{Z}(\mathcal{H}_{3})$ (see Sec. VI.3 for the notations). In this case, we find three stable solutions. Two solutions related by symmetry [i.e. $\mathbf{n}_{1}=\{1,0,0,0,0,0,0,1,2,0,0,0\}^{T}=(\mathbf{1},\pi_{1},2\times\pi_{2})$ and $\mathbf{n}_{1}=(\mathbf{1},\pi_{1},2\times\sigma_{1})$], due to the symmetry between the $\pi_{2}$ and $\sigma_{1}$ quasiparticles, point to the fusion ring $H_{6}$ and one [i.e. $\mathbf{n}_{1}=(\mathbf{1},\pi_{1},\pi_{2},\sigma_{1}$)] to the fusion ring $H_{4}$ [60]. The fusion ring $H_{6}$ corresponds to the UFCs $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$, whereas $H_{4}$ refers to the UFC $\mathcal{H}_{1}$. The two categories $\mathcal{H}_{3}$ and $\mathcal{H}_{2}$ have the same fusion rules but different $F$ symbols.
@@ Appendix D Generalized Hamiltonian
The Hamiltonian that we consider in the main text, Eq. (10), assigns the same energy penalty +1 to each fluxon which is not the vacuum. However, we can assign different energy penalties to each fluxon type. Further, we can also assign a plaquette-dependent energy penalties. More generally, let us define a projector
@@
where $E^{A}_{p}$ is the energy cost of having fluxon $A$ on plaquette $p$ and $\cal F$ is the set of fluxons. So long as $E^{\bf 1}_{p}<E^{A}_{p}$ for all $A\neq{\bf 1}$ and all $p$, then the ground state is still the state where all plaquettes are in the vacuum state. The Hamiltonian we consider in the main text corresponds to $E_{p}^{A}=0$ for all $A\neq{\bf 1}$ and $E_{p}^{\bf 1}=-1$, for all plaquettes $p$.
It is not hard to construct the projectors $B_{p}^{A}$ from the structure of the tube algebra by inserting a projector $P_{A}^{11}$ inside the plaquette $p$ and then fusing it into the edges of the plaquette. For the case where $n_{A,1}>1$, in fact there is freedom to generalize Hamiltonian Eq. (99) further. Since the space of states where a plaquette $p$ has fluxon $A$ is $n_{A,1}$-dimensional, one can add terms to split this added degeneracy. Thus we can even more generally write
@@
Let us take an example of the category Vec($S_{3}$) as discussed in Sec. VI.1. There are three fluxon types labeled $A,B,C$. Here $A$ is the vacuum, which we give energy $E^{A}=0$. Then $B$ we give energy $E^{B}=x$. Finally $C$ is two-dimensional, i.e. $n_{C,1}=2$, and we call $E^{C,1}=y$ and $E^{C,2}=z$ the two values of $E^{C,a}$ with $a=1,2$. For simplicity here, we assume all plaquettes are identical to each other, although this is not necessary.
@@ Appendix E Hilbert-space dimension
Both calculations are actually sums over fusion networks. In Sec. E.1, we are summing over fusion trees of $\cal C$ whereas, in Sec. E.2, we are summing over fusion trees of $\cal Z(\cal C)$ (see the discussion on the derivation of the Moore-Seiberg-Banks formula in Sec. IV.1). In both cases, the fusion trees are sums over fusion multiplicity matrices. In Sec. E.1 we are summing fusion matrices of $\cal C$ whereas in Sec. E.2 we are summing fusion matrices of ${\cal Z}({\cal C})$. These two different fusion multiplicity matrices are related to each other via the commutativity of fusion and restriction [56, 42]. In particular, we have Eq. (97), where $r,s,t\in\cal C$ and $A,B,C\in\cal Z(C)$ and $n_{A,r}$ are lifting coefficients. To clearly distinguish the two sets of fusion matrices we call $\tilde{N}$ that in the input category $\mathcal{C}$ and $N$ that in the output category $\mathcal{Z}(\mathcal{C})$.
@@ E.1 Counting edge labelings
For more complex graphs, one may use restructuring moves (or “$F$-moves”, which here are just the associativity of fusion) to change the connectivity of the graph without altering the Hilbert-space dimension, as shown in Fig. 14.
@@
However, in the case where fusion rules are noncommutative, the situation is different, and on a surface of nonzero genus, one cannot reduce the fusion diagram to a chain of bubbles of the form $L$. To understand this, let us first consider the case $g=1$ (torus). We draw the torus as a rectangle with opposite edges identified. Using the above restructuring moves (Fig. 14), we can reduce any trivalent graph to the following (where the graph edges are black, and we do not draw arrows or labels on these edges for simplicity).
@@
If there are multiple lines running around the handles, Eq. (106) can be reduced to single lines running around the handle by using restructuring moves (Fig. 14). However if there is a single line going around a handle (as shown in the above figure), it cannot be removed by restructuring moves.
@@
where the over-crossing could equally well be drawn as an undercrossing, since it is only the connectivity that matters. But note that due to noncommutativity of the fusion rules, it is not possible to remove the under-crossing and flatten the diagram.
@@
since two vertices occur within the twisted bubble and the remaining vertices are within the planar graph which can be reduced to a bubble chain of $L$’s. Note that for the case of commutative fusion we have $\tilde{L}=L$ and we recover the previous result (105).
To extend this to a genus-$g$ surface, we start with the planar representation of this surface as a $4g$-sided polygon with certain edges identified as in the following picture
@@
Unfortunately, this structure is still somewhat inconvenient, so we use restructuring moves to connect together the legs of each twisted bubble, getting
@@
Now the total Hilbert-space dimension is then